OpenStudy (anonymous):

Find f(5) for this sequence: f(1) = 2 and f(2) = 5, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______

OpenStudy (anonymous):

@jigglypuff314

OpenStudy (anonymous):

ok so f(5) = f(1) + f(2) + f(n-1) and f(n-1) = f(1) + f(2) + f(n-2) ok?

OpenStudy (anonymous):

and 5>2 so we can say f(5) = f(1) + f(2) + f(4) ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok so f(5) = f(1) + f(2) + f(4) what else?

OpenStudy (anonymous):

well and what is f(4)

OpenStudy (anonymous):

oh and f(1) = 2 f(2) = 5 2 + 5 + f(4) f(4) is 12 im guessing

OpenStudy (anonymous):

f(4) = 2 + 5 + f(3) and f(3) = f(1) + f(2) + f(2) = 2 + 5 + 5 = 12 so f(4) = 2 + 5 + 12 = 19

OpenStudy (anonymous):

so f(4) = 19 not 12

OpenStudy (anonymous):

so 2 + 5 + f(4) = ?

OpenStudy (anonymous):

2 + 5 + 19 = 26

OpenStudy (anonymous):

yes :) got it?

OpenStudy (anonymous):

yes, thank you :)

OpenStudy (anonymous):

you're welcome ;)